the hcf of two numbers is 17 and their lcm is 1071.if one of the number is 17 find the other.

Answer by buntibhandari(1)


ans:
dear reader
the formula is 1st no. * 2nd no. = hcf * lcm
and we had hcf and lcm and one no. is given
so just use formula
1st no. * 2nd no. = hcf * lcm
= 17 * 2nd no. = 17*1071
2nd no. = 17*1071/17=1071 ans.